Which sequence of numbers does 2n+3 describe ?

Found 2 solutions by checkley79, Alan3354:
Answer by checkley79(3341)

STARTING @ ZERO THEY ARE:
2*0+3=3
2*3+3=9
2*9+3=21
2*21+3=45

Answer by Alan3354(69443)

Which sequence of numbers does 2n+3 describe ?
------------
Odd numbers, starting at 3.